OpenStudy (thecharmingcharm):

Evaluate 9^P^4 . 15,120 36 3,024 504

alones (alones):

Whoa like this? \(9^P\)\(^{4}\)

OpenStudy (jayrules232):

3,024

OpenStudy (jayrules232):

9 / (9-4) = 362,880 / 120 = 3024

OpenStudy (thecharmingcharm):

Thanks

OpenStudy (mathmate):

Is it possible to evaluate numerically \(9^{P^4}\) without even knowing P? Wow!

OpenStudy (thecharmingcharm):

I meant it like this  Sorry for the confusion :P